OpenStudy (anonymous):

can someone help me or tell me what I am doing wrong please x=-2 -8x=10y+40 -8*2=10y+40 16=10y+40 how do I solve it from here?

OpenStudy (anonymous):

should be -8 * (-2)...not 2

OpenStudy (anonymous):

it should b -8*-2 as above it is x=-2

OpenStudy (anonymous):

right that was a typo...then what? do I subtract 16 from 40 or 40 from 16 and divide? confused

OpenStudy (anonymous):

-8*(-2)=10y+40 16=10y+40 10y=16-40 10y=-24 y=-2.4

OpenStudy (anonymous):

how did you get -2.4

OpenStudy (anonymous):

divide by 10 on both sides...I got it..thank you again...

OpenStudy (anonymous):

Welcome!